OpenStudy (anonymous):

Find a function such that the domain is the set of natural numbers and the output produces the terms of the sequence 8, 12, 16, 20, ... .

OpenStudy (agreene):

y=4(x+1) Check for given values of x: 4(1+1) = 8 4(2+1) = 12 4(3+1) = 16 4(3+1) = 20

OpenStudy (anonymous):

thanks